OpenStudy (anonymous):

Determine which ordered pair is a solution of y = 2x – 3. PLEASE HELP FAST

OpenStudy (anonymous):

THANK YOU !!!

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

wheres your answer it desappeared

OpenStudy (anonymous):

i messed up i'm looking over it

OpenStudy (anonymous):

y = 2x - 3 y = 2*(-3) - 3 y = -6 - 3 y = -9, so plot the point (-3, -9) Now find y when x = 3: y = 2*3 - 3 y = 6 - 3, y = 3, so plot the point (3,3)

OpenStudy (anonymous):

wait but i only have (–2, 7) (0, 3) (–4, –11) (5, –7) as my answers

OpenStudy (anonymous):

? okay hold on lets check

OpenStudy (anonymous):

plug in your values and see which one is true. Process of elimination.

OpenStudy (anonymous):

aint no (0,-3) up there

OpenStudy (anonymous):

y = 2x + 3 Lets try (-2,7) 7 = 2(-2) + 3 7 = -4 + 3 7 = - 1 (not this one) y = 2x + 3.....(0.3) 3 = 2(0) + 3 3 = 0 + 3 3 = 3 (this one works) y = 2x + 3...(-4,-11) -11 = 2(-4) + 3 -11 = - 8 + 3 -11 = - 5 (not this one) y = 2x + 3....(5,-7) -7 = 2(5) + 3 -7 = 10 + 3 -7 = 13 (not this one) So, your answer is (0.3)

OpenStudy (anonymous):

typo....not supposed to be a decimal in the answer....supposed to be (0,3)